Grade 12th passPhysical Chemistry

the crystal of potassium bromide is fa ce centered cubic lattice . its molecular weight and density are 118.91 g/mol and 2.823 g/cm3 respectively then find edge length of its unit cell

shailesh
9 Years agoGrade 12th pass

1 Answer

Vikas TU
9 Years ago
From the density formulae:
2.823 = 4*118.91/Na*V
Volume of unit cell = 28.08*10^-23 cm^3
28.08*10^-23 = a^3
a = 6.542*10^-8 cm